A house is 50.0 ft long and 26 ft wide and has 8.0-ft-high ceilings. What is the volume of the interior of the house in cubic meters and in cubic centimeters?
1
Expert's answer
2019-02-06T12:48:12-0500

We can find the volume of the interior of the house from the formula:

"V = a \cdot b \cdot c,"

here,

"a"

is the length of the house,

"b"

is the width of the house and

"c"

is the height of the house.

Let's first find the volume of the interior of the house in cubic feet:

"V = a \cdot b \cdot c = 50 ft \cdot 26 ft \cdot 8 ft = 10400 ft^3."

Let's convert cubic feet into cubic meters:

"10400 ft^3 \cdot \frac{(0.3048 m)^3}{(1 ft^3 )} = 294.49 m^3."

Let's convert cubic feet into cubic centimeters:

"10400 ft^3 \cdot \frac{(30.48 cm)^3}{(1 ft^3 )} = 294.49 \cdot 10^6 cm^3."

Answer:

"V = 10400 ft^3 = 294.49 m^3 = 294.49 \cdot 10^6 cm^3"

.